The area under a normal curve with mu = 15 and sigma = 2 is
0
1
2
none of these
If Mary gets a 90 on a physics test where the mean is 85 and the standard deviation is 4.59, where does she stand in relation to her classmates?
She is in the top 15%.
She is in the bottom 15%.
She is in the top 5%.
She is in the bottom 5%.
In a normal distribution with mu = 25 and sigma = 9, what number corresponds to z = -2?
21
16
7
none of these

Answer by ewatrrr(10682)

 
Hi,
Note: Bell shaped Curve showing placement of z = 0, z = ±1, 2, 3
Left or Right of z = 0 being 50% of the area under the standard normal curve...
The area under a normal curve = 1

As to: A set of data is normally distributed with a mean of 85 and SD of 4.59
x = 90
z = (90-85)/4.59 = 1.089
Exel function NORMSDIST(1.089) = .862 , 86.2% to the left and 13.8% to the right

As:a normal distribution with mu = 25 and sigma = 9, if z = -2,
then z = (x -25)/9 = -2, and x = -18 + 25 = 7